### v2

@@ Abstract
Tiny object detection is one of the key challenges in the field of object detection. The performance of most generic detectors dramatically decreases in tiny object detection tasks. The main challenge lies in extracting effective features of tiny objects. Existing methods usually perform generation-based feature enhancement, which is seriously affected by spurious textures and artifacts, making it difficult to make the tiny-object-specific features visible and clear for detection. To address this issue, we propose a self-reconstructed tiny object detection (SR-TOD) framework. We for the first time introduce a self-reconstruction mechanism in the detection model, and discover the strong correlation between it and the tiny objects. Specifically, we impose a reconstruction head in-between the neck of a detector, constructing a difference map of the reconstructed image and the input, which shows high sensitivity to tiny objects. This inspires us to enhance the weak representations of tiny objects under the guidance of the difference maps. Thus, improving the visibility of tiny objects for the detectors. Building on this, we further develop a Difference Map Guided Feature Enhancement (DGFE) module to make the tiny feature representation more clear. In addition, we further propose a new multi-instance anti-UAV dataset, which is called DroneSwarms dataset and contains a large number of tiny drones with the smallest average size to date. Extensive experiments on the DroneSwarms dataset and other datasets demonstrate the effectiveness of the proposed method. The code and dataset will be publicly available.
@@ 1 Introduction
The subfield of object detection that identifies and categorizes objects with diminutive dimensions is known as small object detection. According to setting in MS COCO [31], an object qualifies as small if it occupies an area equal to or less than 32×32 pixels. The AI-TOD benchmark [49] refines this definition and categorizes objects as “very tiny” if they span 2 to 8 pixels, “tiny” for 8 to 16 pixels, and “small” for 16 to 32 pixels. We term these objects uniformly as “tiny” objects. Tiny objects commonly appear in a variety of real-world applications, such as autonomous driving, anti-UAV, aerial imagery, industrial inspection, and pedestrian detection, often restricted by considerable imaging distances or the inherently minute size of the objects. Despite its relevance, tiny object detection (TOD) remains an arduous endeavor, with even state-of-the-art detectors struggling to bridge the performance disparity between tiny and normal-sized object detection [11]. The pursuit of improved methods in this area is of considerable theoretical and practical importance.
@@ 2.1 Object Detection
2.2 Tiny Object Detection
Generic detectors have shown good performance in detecting large and medium-sized objects, but they still face significant challenges in detecting tiny objects. In recent years, the research in tiny object detection has mainly focused on data augmentation, scale awareness, context modeling, feature imitation, and label assignment [11].
Data augmentation. Krisantal et al[24] copy and paste small objects to increase the number of samples. DS-GAN [5] designs a new data augmentation pipeline to generate high-quality synthetic data for small objects.
@@ 2.2 Tiny Object Detection
Scale awareness. Lin [29] et al. propose the most popular multi-scale network feature pyramid network using a pyramid of feature levels and feature fusion. Singh et al[44] design the Scale Normalization for Image Pyramids (SNIP) to select some instances for training. PANet [32] enriches the feature hierarchy through bidirectional paths and enhanced deeper features with accurate positioning signals. NAS-FPN [17], Bi-FPN [46] and Recursive-FPN [36] further develop upon the basis of FPN. Gong et al.  [19] adjust the coupling between adjacent layers of FPN by setting fusion factors to optimize feature fusion and improve model performance. Yang et al.  [52]design a cascaded sparse query mechanism to effectively utilize high-resolution features to enhance the detection performance of small objects while maintaining fast inference speed.
Context modeling. Chen et al[9] utilizes context region representations containing proposed patches for subsequent recognition. SINet [22] introduces a context-aware RoI pooling layer to maintain contextual information.
Feature imitation. Many methods [2, 1, 27]utilize generative adversarial networks to perform super-resolution on small objects. Noh et al[35] alleviated the mismatch between the receptive fields of high-resolution features and low-resolution features through dilated convolutions. Deng et al[14] proposed a feature texture transfer module to expand the feature pyramid, enabling the new feature layers to contain more detailed information of small objects.
Label assignment. ATSS [56] adaptively adjusts positive and negative samples based on their statistical features. Xu et al.  [51] propose a simple yet effective strategy called Receptive Field-based Label Assignment (RFLA) to alleviate the scale-sample imbalance problem in anchor-based and anchor-free detectors.
@@
Many small object detection methods have not focused on the key issue of information loss. Feature imitation methods attempt to alleviate this issue through generation but tend to fabricate spurious textures and artifacts [11, 14]. In contrast, we introduce an image self-reconstruction mechanism to identify areas of significant information loss and leverage this prior knowledge to improve small object detection performance.
2.3 Anti-UAV Dataset
@@
Drones, due to their tiny size and low cost among other features, are widely used in fields such as inspection, surveillance, and the military. However, illegal drone flights have introduced many potential risks and threaten public safety, making anti-UAV measures an important task. Since drones are actually tiny in size and often fly at medium to high altitudes, they tend to appear very small in images captured by cameras. Therefore, anti-UAV technology is very well suited for the application of tiny object detection. To our knowledge, there are three publicly available visible light anti-UAV datasets in the current field.
@@
MAV-VID [41]. The dataset consists of 64 video sequences containing single drone captured from multiple viewpoints, with the targets predominantly located in the central region of the images. The drones are relatively large in scale, with an average size of approximately 166 pixels.
Drone-vs-Bird [12]. The dataset consists of 77 video sequences and is used to distinguish between drone and bird targets. The average size of the drones in the dataset is approximately 28 pixels.
DUT Anti-UAV [57].The dataset is divided into two subsets: detection and tracking. The majority of the images in this dataset contain single objects and almost no very tiny objects.
Existing anti-UAV datasets typically contain one drone instance per image alone, with few tiny objects and almost without very tiny objects. This does not match the practical scenarios of anti-UAV measures, which limits the development and application of tiny object detection in anti-UAV tasks. To this end, we introduce the first multi-instance anti-UAV dataset, DroneSwarms, with the smallest average object size to date.
@@ 3.3 Difference Map Guided Feature Enhancement
3.4 DroneSwarms Dataset
Typically, drones operate far from the surveillance apparatus, situated at considerable distances and altitudes, resulting in drone objects that are very tiny and lack clarity. Therefore, the anti-UAV scenario is an important application scenario suitable for tiny object detection. Furthermore, current tiny object detection datasets commonly include many medium and large objects, with average object sizes all above 12.8 pixels [11, 51]. In order to construct a dataset consisting almost entirely of a large number of tiny objects, we propose a object detection dataset with the smallest average size currently for anti-UAV, named DroneSwarms. DroneSwarms consists of 9,109 images and 242,218 annotated UAV instances, with 2,532 used for testing and 6,577 used for training. On average, each image contains 26.59 drone instances. The images are in the size of $1920\times 1080$, manually labeled with high precision. DroneSwarms encompasses a variety of outdoor settings such as urban environments, mountainous terrain, and skies, among others. Unlike existing anti-UAV datasets, DroneSwarms contains 241,249 tiny objects with size of 32 pixels or below, accounting for approximately 99.60%, and the average size is only about 7.9 pixels. The drones are dispersed across the entirety of the image. Therefore, DroneSwarms can be used to comprehensively evaluate methods for tiny object detection .
@@ 4.1 Experimental Setting
Datasets. Experiments are conducted on three datasets. The main experiments are performed on our anti-UAV dataset DroneSwarms, which has the smallest average absolute object size of about 7.9 pixels. Besides, we conduct experiments on two other datasets: VisDrone2019 [15] and AI-TOD [49] datasets. The VisDrone2019 [15] encompasses objects spanning 10 categories, and all images are captured from the perspective of drones. AI-TOD [49] comprises objects across 8 categories, and the images are collected from various datasets containing tiny objects. All selected datasets contain a large number of tiny objects, especially tiny objects with sizes below 16 pixels.
Implementation Details. All core codes are built upon MMdetection [10]. Since the DroneSwarms dataset mainly consists of a large number of tiny objects, in order to ensure model convergence, we set the initial learning rate to 0.0025 for all experiments on the DroneSwarms dataset, and use the Stochastic Gradient Descent (SGD) optimizer training models for 20 epochs with 0.9 momenta, 0.0001 weight decay, 2 batch size, and 2 anchor scale. For experiments on VisDrone2019 [15] and AI-TOD [49], we strictly followed all the experiment settings of RFLA [51] on these two datasets, such as using the SGD optimizer for 12 epochs training. And the initial learning rate is set to 0.005 and decays at the 8th and 11th epochs. The ImageNet [43] pre-trained model, ResNet-50, is used as the backbone. All the other parameters of baselines are set the same as default in MMdetection. We used 2 NVIDIA RTX 3090 GPUs for model training on DroneSwarms and 1 NVIDIA RTX 3090 GPU for model training on the other two datasets. The above parameters are used in all experiments unless specified otherwise.
Evaluation Metrics. While the Average Precision (AP) evaluation metric from MS COCO [31] is prevalent among object detection algorithms, it is developed only with generic detectors in mind. This metric broadly categorizes objects with size under 32 pixels as small, employing only the APs for their assessment. Thus, to more effectively demonstrate our method’s performance in detecting small objects of diverse sizes, the evaluation metric follows AI-TOD benchmark [49]. Note that APvt,APt, APs are APs for very tiny, tiny, small scales, respectively. The definitions of these scales are introduced in Sec. 1.
4.2 Results on DroneSwarms
@@
We compare our method with other state-of-the-art detectors on DroneSwarms dataset, as shown in Tab. 1. RFLA w/ SR-TOD achieves an AP of 39.0, surpassing state-of-the-art competitors by 1.1 AP and resulting in an overall performance boost of 2.1 AP. Notably, RFLA w/ SR-TOD registers a substantial performance gain of 2.3 points in both APvt and APt, underscoring the pronounced efficacy of our method in detecting tiny objects. Furthermore, within the Cascade R-CNN framework, Cascade R-CNN with SR-TOD exceeds RFLA by 1.4 AP, indicating a notable improvement of 1.9 AP. Moreover, our method has also manifested clear performance gains when applied to other detectors. One-stage detectors typically show lower performance compared to multi-stage detectors due to the lack of multi-stage regression [54, 59, 15, 51]. Even though the one-stage detector RetinaNet [30] only utilizes the feature map $P3$ with lower resolution from FPN, our method achieves a 1.4 AP improvement. It is particularly noteworthy that, due to focusing on different challenges of tiny object detection, our framework can work in tandem with other methods like RFLA. And our framework does not conflict with methods like DetectoRS [36], which also enhance feature representation by improving FPN. Compared to the baseline methods, our method achieves consistent improvements.
4.3 Results on VisDrone2019 and AI-TOD
@@
Table 3: Results on AI-TOD. The best results are highlighted.
@@
We further verify the effectiveness of DGFE on VisDrone2019 [15] and AI-TOD [49], which simultaneously hold a large scale variance and contain many tiny objects of different categories, as shown in Tab. 2 and Tab. 3.
@@
AI-TOD. DetectoRS w/ SR-TOD attains an AP of 24 on the AI-TOD dataset [49], exceeding RFLA by a margin of 2.3 AP and significantly surpassing all competitors. All detectors have significant performance improvements. Significantly, Cascade R-CNN w/ SR-TOD surpasses RFLA by 1.3 points in APvt and by 0.6 points in APt, underscoring the substantial impact of our method on tiny object detection within the AI-TOD dataset [49].
The consistent improvement on various datasets indicates the generality of SR-TOD.
@@ 4.4 Ablation Study and Discussion
Effectiveness of individual component. As shown in Sec.3.2 and Sec.3.3, the self-reconstruction mechanism we propose consists of two key components: the Reconstruction Head and the Difference Map Guided Feature Enhancement module. Note that the validation of DGFE relies on the Reconstruction Head, as it requires the construction of a difference map as input. When only the reconstruction head is applied, we obtain the reconstructed image without constructing a difference map. We progressively apply Reconstruction Head and DGFE into the Cascade R-CNN. Results are listed in Tab. 4.3. The application of DGFE resulted in 1.8 points increase in AP over the use of the reconstruction head alone, accompanied by a substantial gain of 2.1 points in APvt. These results conclusively prove that DGFE is both an effective and indispensable module for leveraging the difference map. It should be highlighted that the sole application of the reconstruction head yielded a modest improvement of 0.1 AP, along with an increment of 0.2 points in both APt and APs. This enhancement can be attributed to the fact that integrating image reconstruction constraints within the object detection model not only preserves detection performance but also, to a certain degree, improves the backbone network’s pixel comprehension. This approach could potentially offer fresh inspiration for the field of object detection.
@@
Effectiveness of threshold filtration. As shown in Sec. 3.3, we implement an operation within the DGFE that filters the difference map using learnable thresholds. To examine the impact of thresholding, we also explore methods employing fixed thresholds as well as those devoid of any threshold-based filtering. To more robustly assess the role of thresholds, we conduct experiments on the VisDrone2019 dataset [15], which features intricate backgrounds and diverse categories. The results are shown in Tab. 4.3. Undoubtedly, the approaches that involve setting a threshold generally exceed the performance of the method without a threshold, thus affirming the utility of thresholding. In addition, the learnable threshold method performs similarly to the fixed threshold method in APvt but shows a 0.3 points improvement in APt and a 0.7 points improvement in APs. This indicates that tiny objects with size above 8 pixels are more sensitive to thresholds. Therefore, the method with learnable threshold achieves better performance by finely adjusting the threshold.
@@
Exploration of high-frequency difference map. Recently, some image reconstruction methods [58, 23] model images in the frequency domain. The high-frequency information of an image includes details such as edges and textures, which are precisely the parts that are difficult to recover in image self-reconstruction. In particular, noise is also considered part of high-frequency information, and very tiny objects can be approximated as noise-like points with a size of only a few pixels. Inspired by this, we further utilize the fast Fourier transform algorithm to separate the high-frequency components of the original input image and the reconstructed image, and construct high-frequency difference maps, as shown in the Fig. 3. It can be clearly seen that although the high-frequency difference map and pixel difference map are quite similar, the high-frequency difference map has more refined contours on some drone objects. Additionally, the high-frequency difference map significantly reduces noise signals. However, this also results in certain smaller drone objects appearing even more faint in the high-frequency difference map. Then we replace the pixel difference map (PD) with the high-frequency difference map (HFD) and conduct experiments on DroneSwarms using Cascade R-CNN w/ SR-TOD. The results are shown in Tab. 4.3. It can be seen that HFD has a 0.1 AP improvement compared to PD. Moreover, there is an improvement of 0.3 points in APvt and 0.2 points in APt. This demonstrates that high-frequency difference map can provide more accurate priors for tiny objects compared to pixel difference map. Therefore, we think that the information severely lost in tiny object feature extraction mainly lies in the high-frequency components. Note that compared to the baseline, the detection performance of using pixel difference map and high-frequency difference map is very similar. Considering computational efficiency, we use pixel difference map as the basic setting for SR-TOD. However, the high-frequency difference map shows the potential for further exploration of our method. More ablation experiments will be provided in the supplementary materials.
Figure 4: Visualizations on DroneSwarms. The first row is the local regions of input images. The second row is the corresponding feature maps from Cascade R-CNN. The third row is the feature maps from Cascade R-CNN w/ SR-TOD.

### v3

@@ Abstract
Tiny object detection is one of the key challenges for most generic detectors. The main difficulty lies in extracting effective features of tiny objects. Existing methods usually perform generation-based feature enhancement, which is seriously affected by spurious textures and artifacts, making it difficult to make the tiny-object-specific features visible and clear for detection. To address this issue, we propose a self-reconstructed tiny object detection (SR-TOD) framework. We for the first time introduce a self-reconstruction mechanism in the detection model, and discover the strong correlation between it and the tiny objects. Specifically, we impose a reconstruction head in-between the neck of a detector, constructing a difference map of the reconstructed image and the input, which shows high sensitivity to tiny objects. This inspires us to enhance the weak representations of tiny objects under the guidance of the difference maps. Thus, improving the visibility of tiny objects for the detectors. Building on this, we further develop a Difference Map Guided Feature Enhancement (DGFE) module to make the tiny feature representation more clear. In addition, we further propose a new multi-instance anti-UAV dataset. Extensive experiments demonstrate our effectiveness. The code is available: https://github.com/Hiyuur/SR-TOD.
@@ 1 Introduction
The subfield of object detection that identifies and categorizes objects with diminutive dimensions is known as tiny object detection. According to setting in MS COCO [32], an object qualifies as small if it occupies an area equal to or less than 32×32 pixels. The AI-TOD benchmark [51] refines this definition and categorizes objects as “very tiny” if they span 2 to 8 pixels, “tiny” for 8 to 16 pixels, and “small” for 16 to 32 pixels. We term these objects uniformly as “tiny” objects. Tiny objects commonly appear in a variety of real-world applications, such as autonomous driving, industrial inspection, and pedestrian detection, often restricted by considerable imaging distances or the inherently minute size of the objects. Despite its relevance, tiny object detection (TOD) remains an arduous endeavor, with even state-of-the-art detectors struggling to bridge the performance disparity between tiny and normal-sized object detection [11]. The pursuit of improved methods in this area is of considerable theoretical and practical importance.
@@ 2 Related Work
Tiny Object Detection is a challenging task for most generic detectors. In recent years, the research in tiny object detection has mainly focused on data augmentation, scale awareness, context modeling, feature imitation, and label assignment [11]. Data augmentation: Krisantal et al[25] copy and paste tiny objects to increase the number of samples. DS-GAN [5] designs a new data augmentation pipeline to generate high-quality synthetic data for tiny objects. Scale awareness: Lin [30] et al. propose the most popular multi-scale network feature pyramid network using a pyramid of feature levels and feature fusion. Singh et al[46] design the Scale Normalization for Image Pyramids (SNIP) to select some instances for training. PANet [33] enriches the feature hierarchy through bidirectional paths and enhanced deeper features with accurate positioning signals. NAS-FPN [17], Bi-FPN [48] and Recursive-FPN [38] further develop upon the basis of FPN. Gong et al.  [19] adjust the coupling between adjacent layers of FPN by setting fusion factors to optimize feature fusion. Yang et al.  [54] design a cascaded sparse query mechanism to effectively utilize high-resolution features to enhance the detection performance of tiny objects while maintaining fast inference speed. Context modeling: Chen et al[9] utilizes context region representations containing proposed patches for subsequent recognition. SINet [23] introduces a context-aware RoI pooling layer to maintain contextual information. Feature imitation: Many methods [2, 1, 28]utilize generative adversarial networks to perform super-resolution on tiny objects. Noh et al[37] alleviated the mismatch between the receptive fields of high-resolution features and low-resolution features through dilated convolutions. Deng et al[14] proposed a feature texture transfer module to expand the feature pyramid, enabling the new feature layers to contain more detailed information of tiny objects. Label assignment: ATSS [60] adaptively adjusts positive and negative samples based on their statistical features. Xu et al.  [53] propose a simple yet effective strategy called Receptive Field-based Label Assignment (RFLA) to alleviate the scale-sample imbalance problem in anchor-based and anchor-free detectors.
@@
Many tiny object detection methods have not focused on the key issue of information loss. Feature imitation methods attempt to alleviate this issue through generation but tend to fabricate spurious textures and artifacts [11, 14]. In contrast, we introduce an image self-reconstruction mechanism to identify areas of significant information loss and leverage this prior knowledge to improve tiny object detection performance.
@@
Anti-UAV Dataset. Drones, due to their tiny size and low cost among other features, are widely used in inspection and surveillance. However, illegal drone flights have introduced many potential risks and threaten public safety, making anti-UAV measures an important task. Since drones are tiny in size and often fly at medium to high altitudes, they tend to appear very small in images. Therefore, anti-UAV technology is very well suited for the application of tiny object detection.
@@
To our knowledge, there are three publicly available visible light anti-UAV datasets in the current field. MAV-VID [43] consists of 64 video sequences containing a single drone captured from multiple viewpoints, with the targets predominantly located in the central region of the images. The drones are relatively large in scale, with an average size of approximately 166 pixels. Drone-vs-Bird [12] consists of 77 video sequences and is used to distinguish between drone and bird targets. The average size of the drones in the dataset is approximately 28 pixels. DUT Anti-UAV [61] is divided into two subsets: detection and tracking. The majority of the images in this dataset contain single objects and almost no very tiny objects.
@@ 4.1 Experimental Setting
Datasets. The experiments are conducted on three datasets. The VisDrone2019 [15] encompasses objects spanning 10 categories, and all images are captured from the perspective of drones. The AI-TOD dataset [51] comprises objects across 8 categories, and the images are collected from various datasets containing tiny objects. Our anti-UAV dataset DroneSwarms has the smallest average absolute object size of about 7.9 pixels. All selected datasets contain a large number of tiny objects, especially tiny objects with sizes below 16 pixels.
@@
Implementation Details. All core codes are built upon MMdetection [10]. Since the DroneSwarms dataset mainly consists of a large number of tiny objects, in order to ensure model convergence, we set the initial learning rate to 0.0025 for all experiments on the DroneSwarms dataset, and use the Stochastic Gradient Descent (SGD) optimizer training models for 20 epochs with 0.9 momenta, 0.0001 weight decay, 2 batch size, and 2 anchor scale. For experiments on VisDrone2019 [15] and AI-TOD [51], we strictly followed all the experiment settings of RFLA [53] on these two datasets, such as using the SGD optimizer for 12 epochs training. And the initial learning rate is set to 0.005 and decays at the 8th and 11th epochs. The ImageNet [45] pre-trained model, ResNet-50, is used as the backbone. All the other parameters of baselines are set the same as default in MMdetection. Our experiments are conducted on Huawei Atlas 800 Training Server with CANN and NVIDIA RTX 3090 GPU.
Evaluation Metrics. While the Average Precision (AP) evaluation metric from MS COCO [32] is prevalent among object detection algorithms, it is developed only with generic detectors in mind. This metric broadly categorizes objects with size under 32 pixels as small, employing only the APs for their assessment. Thus, to more effectively demonstrate our method’s performance in detecting tiny objects of diverse sizes, the evaluation metric follows AI-TOD benchmark [51]. Note that APvt,APt, APs are APs for very tiny, tiny, small scales, respectively. The definitions of these scales are introduced in Sec. 1.
@@
4.2 Results on DroneSwarms
As shown in Tab. 1, RFLA w/ SR-TOD achieves an AP of 39.0, surpassing the competing methods by 1.1 AP and resulting in an overall performance boost of 2.1 AP. Notably, RFLA w/ SR-TOD registers a substantial performance gain of 2.3 points in both APvt and APt, underscoring the pronounced efficacy of our method in detecting tiny objects. Furthermore, within the Cascade R-CNN framework, Cascade R-CNN with SR-TOD exceeds RFLA by 1.4 AP, indicating a notable improvement of 1.9 AP. Moreover, our method has also manifested clear performance gains when applied to other detectors. One-stage detectors typically perform worse than multi-stage detectors due to the lack of multi-stage regression [56, 63, 15, 53]. Even though the one-stage detector RetinaNet [31] only utilizes the feature map $P3$ with lower resolution from FPN, our method achieves a 1.4 AP improvement. It is particularly noteworthy that, due to focusing on different challenges of tiny object detection, our framework is qualified to work with other methods like RFLA. Our framework does not conflict with methods like DetectoRS [38], which also enhance feature representation by improving FPN. Compared to the baseline methods, our method achieves consistent improvements. We have also conducted experiments with transformer-based detectors on the DroneSwarms dataset. Due to the slow convergence speed of transformer-based detectors, the schedule for Deformable-DETR [64] is set to 50 epochs as recommended. Obviously, our method shows significant superiority, even compared with the advanced transformer-based detector DINO [59]. Besides, we have also integrated our SR-TOD with DINO, which introduces improvements of 0.2 AP and 0.9 APs. This further indicates our effectiveness in finding tiny objects when integrating with different architectures, including transformer and CNN. In addition, these experiments that transformer-based detectors shows limited performance in detecting tiny objects. Recent studies also indicated that DETR lacks effective multi-scale features and that the sparse query paradigm does not adequately cover tiny objects [8, 11]. Moreover, the cost of using high-resolution feature maps beneficial for tiny objects in Transformer-based detectors is prohibitively high due to attention mechanisms [64].
4.3 Results on VisDrone2019 and AI-TOD
Table 3: Results on AI-TOD. The best and second results are highlighted.
@@ 4.3 Results on VisDrone2019 and AI-TOD
We further compare our model with the competing methods on VisDrone2019 [15] and AI-TOD [51], as shown in Tab. 2 and Tab. 3.
@@
AI-TOD. DetectoRS w/ SR-TOD attains an AP of 24 on the AI-TOD dataset [51], exceeding RFLA by a margin of 2.3 AP and significantly surpassing all competitors. All detectors have significant performance improvements. Significantly, Cascade R-CNN w/ SR-TOD surpasses RFLA by 1.3 points in APvt and by 0.6 points in APt, underscoring the substantial impact of our method on tiny object detection within the AI-TOD dataset [51]. Our method outperforms HANet [21] on most metrics (e.g., AP: $22.1\rightarrow 24.0$, APt: $22.2\rightarrow 24.8$), indicating our superiority. HANet [21] aims to obtain scale-specific feature subspaces with a predicted activation map. Although the shallow features further compensate for the detailed information of the specific scale feature space, it is still hard to capture tiny objects as their weak signals are not enhanced in shallow layers. Compared to the activation map of HANet, our self-reconstructed difference map is more sensitive to information loss, especially for tiny objects. The consistent improvement on various datasets indicates the generality of SR-TOD. Experiments on more datasets are provided in the supplementary materials.
@@ 4.4 Ablation Study and Discussion
Effectiveness of individual component. As shown in Sec.3.2 and Sec.3.3, the self-reconstruction mechanism consists of two key components: the Reconstruction Head and the Difference Map Guided Feature Enhancement module. Note that the validation of DGFE relies on the Reconstruction Head, as it requires the construction of a difference map as input. When the reconstruction head is applied alone, we obtain the reconstructed image without constructing a difference map. We add Reconstruction Head and DGFE into the Cascade R-CNN one-by-one. The results are listed in Tab. 4.3. The DGFE increased 1.8 points in AP over the use of the reconstruction head alone, accompanied by a substantial gain of 2.1 points in APvt. These results conclusively prove that DGFE is both an effective and indispensable module for leveraging the difference map. It should be highlighted that the sole application of the reconstruction head yielded a modest improvement of 0.1 AP, along with an increment of 0.2 points in both APt and APs. This enhancement can be attributed to the fact that integrating image reconstruction constraints within the object detection model not only preserves detection performance but also, to a certain degree, improves the backbone network’s pixel comprehension. This approach could potentially offer fresh inspiration for the field of object detection.
Figure 4: Visualizations on DroneSwarms. The first row is the local regions of input images. The second and third rows are the corresponding feature maps from Cascade R-CNN $w/o$ and $w/$ SR-TOD.
@@
Effectiveness of threshold filtration. To investigate the effect of thresholds, we also explore methods employing fixed thresholds and those without threshold-based filtering. We conduct experiments on the VisDrone2019 dataset [15], which features intricate backgrounds and diverse categories. The results are shown in Tab. 4.3. Undoubtedly, the approaches that involve setting a threshold generally exceed the performance of the method without a threshold, thus affirming the utility of thresholding. In addition, the learnable threshold method performs similarly to the fixed threshold method in APvt but shows a 0.3 points improvement in APt and a 0.7 points improvement in APs. This indicates that tiny objects with a size above 8 pixels are more sensitive to thresholds. Therefore, the method with a learnable threshold achieves better performance by finely adjusting the threshold.
@@
Exploration of the high-frequency difference map. Recently, some image reconstruction methods [62, 24] model images in the frequency domain. The high-frequency information of an image includes details such as edges and textures, which are precisely the parts that are difficult to recover in image self-reconstruction. In particular, noise is also considered part of high-frequency information, and very tiny objects can be approximated as noise-like points with a size of only a few pixels. Inspired by this, we further utilize the fast Fourier transform algorithm to separate the high-frequency components of the original input image and the reconstructed image, and construct high-frequency difference maps, as shown in Fig. 3. Although the high-frequency difference map and pixel difference map are quite similar, the high-frequency difference map has more refined contours on some drone objects. Additionally, the high-frequency difference map significantly reduces noise signals. However, this also results in certain smaller drone objects appearing even more faint in the high-frequency difference map. Then we replace the pixel difference map (PD) with the high-frequency difference map (HFD) and conduct experiments on DroneSwarms using Cascade R-CNN w/ SR-TOD. The results are shown in Tab. 4.3. The HFD has a 0.1 AP improvement compared to PD. Moreover, there is an improvement of 0.3 points in APvt and 0.2 points in APt. This demonstrates that the high-frequency difference map can provide more accurate priors for tiny objects compared to the pixel difference map. This indicates that the information severely lost in tiny object feature extraction mainly lies in the high-frequency components. Note that compared to the baseline, the detection performance of using pixel difference map and high-frequency difference map is very similar. Considering computational efficiency, we use the pixel difference map as the basic setting for SR-TOD. However, the high-frequency difference map shows the potential for further exploration of our method. More results are provided in the supplementary materials.